If, pq+qr+rp= 0 then, (p^2)/(p^2-qr)+ (q^2)/(q^2-rp)+( r^2)/(r^2-pq) =?
1
Expert's answer
2020-06-08T19:48:04-0400

Question:

If, pq+qr+rp= 0 then, (p^2)/(p^2-qr)+ (q^2)/(q^2-rp)+( r^2)/(r^2-pq) =?

Solution:

1 - step: pq = -(qr + rp), qr = -(rp + qp), rp = -(pq + qr).

2 - step: p2/(p2+rp+qp)+q2/(q2+qp+qr)+r2/(r2+qr+rp).

3 - step: p/(p+r+q)+q/(q+p+r)+r/(r+q+p) = (r+q+p)/(r+q+p)=1

Answer: 1.